A ball is thrown horizontally from a height of 17.1 m and hits the ground with a speed that is three times its initial speed. What was the initial speed?
Answer #1

from kinematic equation

v2 - u2 = 2as
(3u)2 - u2 = 2 * 9.8 *17.1

2 u2 =335.16

u=12.94 m/s